# Separable models for dynamic signed networks

### v1

@@ Abstract
Signed networks capture the polarity of relationships between nodes, providing valuable insights into complex systems where both supportive and antagonistic interactions play a critical role in shaping the network’s dynamics. We propose a separable temporal generative framework based on multi-layer exponential random graph models, characterised by the assumption of conditional independence between the sign and interaction effects. This structure preserves the flexibly and explanatory power inherent in the binary network specification while adhering to consistent balance theory assumptions. Using a fully probabilistic Bayesian paradigm, we infer the doubly intractable posterior distribution of model parameters via an adaptive Metropolis-Hastings approximate exchange algorithm. We illustrate the interpretability of our model by analysing signed relations among U.S. Senators during Ronald Reagan’s second term (1985–1989). Specifically, we aim to understand whether these relations are consistent and balanced or reflect patterns of supportive or antagonistic alliances.
Dynamic networks,
Exponential random graph models,
Bayesian inference,
@@
[…]
and
@@ 1 Introduction
A central question in the analysis of signed networks is how well positive and negative edges align with Structural Balance Theory (SBT) (Cartwright and Harary, 1956). SBT asserts that positive relations should exhibit transitivity, as expressed by the saying “the friend of my friend is my friend,” while negative relations should demonstrate anti-transitivity, as in “the enemy of my enemy is my friend.”
@@
To explore these dynamics, our approach utilises the exponential random graph modelling (ERGM) framework (Lusher, Koskinen and Robins, 2013), which is particularly suited to modelling the complex interplay between positive and negative edge configurations. ERGMs allow us to incorporate terms that capture relational tendencies, such as the clustering of positive edges or the formation of unexpected negative edges between individuals with similar nodal covariate information. By accounting for these relational patterns, we can accurately assess the statistical significance of balanced configurations and their impact on the overall structure of the network (Fritz et al., 2022).
In this paper, we examine undirected signed network panel data, which consists of cross-sectional snapshots of network relationships observed at regular intervals. In particular, we focus on the analysis of political relationships among U.S. Senators, measured across successive U.S. Congresses (Neal, 2020, 2014). Each snapshot captures the state of signed interactions within the Senate during each legislative year, offering insight into the evolving dynamics of political alliances and rivalries over time.
@@
We extend the temporal exponential random graph model (TERGM) framework (Robins and Pattison, 2001; Hanneke, Fu and Xing, 2010) to better accommodate the dynamics of signed interactions in longitudinal network data. Specifically, we incorporate a separable specification (Krivitsky and Handcock, 2014) that distinguishes between the formation and persistence of edges, enabling a more detailed understanding of temporal dependence. Building on the decomposition introduced by Lerner (2016), we further enhance the TERGM by modelling the signed network as the outcome of two interrelated processes: an interaction process and a conditional sign process. This extension allows for a more flexible and interpretable parameterisation that captures not only how relationships form and persist over time, but also how structural balance – through both balanced and imbalanced signed configurations – shapes the evolution of network interactions.
@@
The paper is structured as follows. In Sections 2 and 3, we provide an overview of ERGMs and temporal ERGMs respectively. Section 4 offers a comprehensive definition of signed networks and presents an overview of Structural Balance Theory (SBT). In Sections 5 and 6, we introduce a multi-layer interpretation of the signed process, distinguishing between the interaction process and the conditional sign process. We describe how these two components are incorporated into the ERGM generative framework to capture the dynamics of signed relationships. Section 7 combines temporal separability with layer separability to define separable 2-layer temporal ERGMs. Here, we define a Markov process based on binary ERGM terms, which allows for a flexible handling of Structural Balance Theory configurations. In Section 8, we outline a Bayesian inferential approach, which quantifies the uncertainty in the posterior model parameter estimates by using an adaptive Metropolis-Hastings approximate exchange algorithm (Haario, Saksman and Tamminen, 2001; Caimo and Friel, 2011) to facilitate sampling from the doubly-intractable posterior distribution. Finally, Section 9 demonstrates the utility of our modelling approach by applying it to the analysis of political relationships among U.S. Senators. We assess model fit using posterior predictive checks to evaluate the model performance. Section 10 concludes with a discussion on potential extensions and future directions.
@@ 2 Exponential random graph models
where $\mathbf{s}(\mathbf{y})\in\mathbb{R}^{Q}_{>0}$ is a known vector of $Q$ sufficient statistics (Besag, 1974), $\boldsymbol{\varphi}\in\mathbb{R}^{Q}$ is the associated parameter vector, and $\kappa(\boldsymbol{\varphi})$ a normalising constant which is computationally intractable to evaluate for all but trivially small graphs. The dependence hypothesis at the basis of these models is that edges self organize into small structures called configurations. There is a wide range of possible network configurations (Robins et al., 2007) which give flexibility to adapt to different contexts. A positive (or negative) value of the parameter $\varphi$ indicates a tendency for the configuration represented by $s(\mathbf{y})$ to appear more (or less) frequently in the data than would be expected under an Erdős–Rényi random graph model with an edge probability of 0.5.
@@ 3 Temporal exponential random graph models
Temporal exponential random graph models (TERGMs) (Robins and Pattison, 2001; Hanneke, Fu and Xing, 2010) describe the joint distribution of a dynamic network sequence by assuming:
$p(\mathbf{y}_{1},\dots,\mathbf{y}_{T}\mid\mathbf{y}_{0},\boldsymbol{\varphi})$ $=\prod_{t=1}^{T}p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\dots,\mathbf{y}_{0},\boldsymbol{\varphi}\right)$
$=\prod_{t=1}^{T}\exp\{\boldsymbol{\varphi}^{\top}\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1},\dots,\mathbf{y}_{0}\right)-\kappa(\boldsymbol{\varphi},\mathbf{y}_{t-1}\dots,\mathbf{y}_{0})\},$
where $\boldsymbol{\varphi}$ parametrises the influence of sufficient network statistics $\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1},\dots,\mathbf{y}_{0}\right)$ on the likelihood and $\kappa(\boldsymbol{\varphi},\mathbf{y}_{t-1},\dots,\mathbf{y}_{0})$ is a normalising constant.
Assuming first-order Markov dependence simplifies the model by limiting the dependence of $\mathbf{y}_{t}$ to $\mathbf{y}_{t-1}.$ This is of course computationally convenient and empirically justified in many settings:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\dots,\mathbf{y}_{0},\boldsymbol{\varphi}\right)=p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}\right)\propto\exp\left\{\boldsymbol{\varphi}^{\top}\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1}\right)\right\}.$ (2)
@@
where $\boldsymbol{\varphi}^{\mathcal{F}}$ and $\boldsymbol{\varphi}^{\mathcal{P}}$ describe respectively the edge formation and persistence process from $t-1$ to $t.$ Then, given $\mathbf{y}^{\mathcal{F}},\mathbf{y}^{\mathcal{P}}$ and $\mathbf{y}_{t-1},$ we have $\mathbf{y}_{t}=\mathbf{y}^{\mathcal{P}}\cup\left(\mathbf{y}^{\mathcal{F}}\;\backslash\;\mathbf{y}_{t-1}\right).$
The STERGM framework addresses a key limitation of standard TERGMs by separating the modelling of edge formation and persistence. This separation allows for more interpretable and flexible modelling of dynamic networks, especially when edge formation and dissolution are governed by different mechanisms. For example, in a friendship network, a new friendship (formation) and the maintenance of an existing friendship (persistence) may be influenced by different factors — proximity may drive formation, while mutual friends may support persistence.
@@ 4 Signed network structures
A signed network consists of a set of $n$ nodes and signed edges between pairs of them. Its relational structure is described by a random adjacency matrix with elements the set of edge variables $y_{ij}\in\{+1,-1,0\}$ respectively indicating a positive, negative of no relation between node $i$ and node $j.$ As for binary networks, signed networks may be directed or undirected depending on the nature of the relationships between the nodes. In this paper, we focus on undirected signed networks, although the methodologies introduced can be easily extended to the directed case.
Structural balance theory (SBT) examines how positive and negative relationships between actors evolve and ultimately stabilise within a network. It focuses on the dynamics and eventual equilibrium of these relationships in a signed network (Cartwright and Harary, 1956). SBT is commonly illustrated using the familiar concepts of friendship and enmity. These terms provide a clear and intuitive understanding of the theory’s fundamental principles. Balanced triadic configurations include: (1) a friend of a friend is a friend, and (2) an enemy of a friend is an enemy. In contrast, unbalanced configurations are: (3) a friend of an enemy is an enemy, and (4) an enemy of an enemy is a friend. Configuration (1) corresponds to positive triangles (Figure 1 (a)), while configuration (2) corresponds to triangles with two negative edges (Figure 1 (b)). These two patterns are part of the so-called strong balance structure (Harary, 1953). Davis (1967) later introduced the weak balance structure, under which configurations comprising triangles with all negative edges (Figure 1 (c)) are also regarded as balanced. Conversely, triadic configurations with exactly two positive edges and one negative edge (Figure 1 (d)) are considered unbalanced.
@@ 5 Multi-layer interpretation
Multi-layer ERGMs (Krivitsky, Koehly and Marcum, 2020) extend the traditional ERGM framework to model networks with multiple types of interactions between nodes, with each layer representing a distinct kind of relationship.
In the context of signed networks, we can decompose the signed edge process into the product of two network layers $\mathbf{y}=\mathbf{x}\mathbf{z},$ where $\mathbf{x},$ with dyadic elements $x_{ij}=0,$ if $i\sim j$; $x_{ij}=1,$ if $i\not\sim j,$ is a binary adjacency matrix encoding the presence of edges between nodes; and $\mathbf{z}$ with dyadic elements $z_{ij}=-1,$ if $i\overset{-}{\sim}j$; $z_{ij}=+1,$ if $i\overset{+}{\sim}j$ is a binary adjacency matrix encoding the sign of the edges between nodes. Clearly, this can be viewed as a special case of a multi-layer network, where the sign information is meaningful only in the presence of an actual interaction. In other words, the sign layer is effectively constrained by the interaction layer. This dependency will be further discussed in the following section.
@@
where $\mathbf{s}(\mathbf{x},\mathbf{z})$ incorporates network statistics that model the two layers jointly. Figure 2 (a). The modelling approach for signed networks proposed by Fritz et al. (2022) is equivalent to a 2-layer ERGM which represents an ERGM with a uniform categorical reference measure.
@@ 6 Separable modelling
Lerner (2016) highlights the difference between marginal and conditional tests of structural balance, showing that other network effects affecting the likelihood of interaction can hide structural balance effects. This can make it seem as though balance theory is incorrect, even when nodes actually prefer balanced triangles over unbalanced ones. For this reason, to properly test balance theory, we should specify and estimate the conditional probability of an edge being positive or negative.
Unlike the joint model described in the previous section, this separable approach allows for conditional independent modelling of the structure of edges and their signs.
This has practical benefits: it simplifies model estimation, and it also aligns with situations where the factors driving connections (like shared interests or affiliations) are distinct from those affecting relationship quality (like competition or past conflict). For example, in a social network context, the presence of an edge between nodes might be explained by organisational requirements or physical proximity, while the sign could depend on trust, cooperation, or conflict.
The marginal probability of $\mathbf{y}$ can be decomposed into the product of the interaction process $\mathbf{x}$ and the conditional sign process given the presence of interaction $\mathbf{z}\mid\mathbf{x}$:
@@
Figure 2: Graphical representation of (a) 2-layer ERGMs defined in Equation (4) and (b) separable 2-layer ERGMs defined in Equation (5).
7 Separable 2-layer temporal models
The separable model defined in Equation (5), when combined with the separable temporal model family specified in Equation (3), results in a separable 2-layer TERGM, with the transition probability defined as:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\right)=$ $\;p\left(\mathbf{x}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\right)\times p\left(\mathbf{z}_{t}\mid\mathbf{x}_{t},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\right)$ (6)
$=$ $\;p\left(\mathbf{x}_{t}^{\mathcal{F}}\mid\mathbf{x}_{t-1},\boldsymbol{\xi}^{\mathcal{F}}\right)\times p\left(\mathbf{x}_{t}^{\mathcal{P}}\mid\mathbf{x}_{t-1},\boldsymbol{\xi}^{\mathcal{P}}\right)\times$
$\;p\left(\mathbf{z}_{t}^{\mathcal{F}}\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}}\right)\times p\left(\mathbf{z}_{t}^{\mathcal{P}}\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{P}}\right)$
$=$ $\exp\left\{{\boldsymbol{\xi}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{F}}};\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\zeta}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{F}}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\xi}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{P}}};\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\zeta}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{P}}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1})\right\},$

### v2

@@ Abstract
Signed networks capture the polarity of relationships between nodes, providing valuable insights into complex systems where both supportive and antagonistic interactions play a critical role in shaping the network dynamics. We propose a separable temporal generative framework based on multi-layer exponential random graph models, characterised by the assumption of conditional independence between the sign and interaction effects. This structure preserves the flexibly and explanatory power inherent in the binary network specification while adhering to consistent balance theory assumptions. Using a fully probabilistic Bayesian paradigm, we infer the doubly intractable posterior distribution of model parameters via an adaptive Metropolis-Hastings approximate exchange algorithm. We illustrate the interpretability of our model by analysing signed relations among U.S. Senators during Ronald Reagan’s second term (1985–1989). Specifically, we aim to understand whether these relations are consistent and balanced or reflect patterns of supportive or antagonistic alliances.
@@
Dynamic signed networks, structural balance, exponential random graph models, Bayesian inference
@@ 1 Introduction
A central question in the analysis of signed networks is how well positive and negative edges align with Structural Balance Theory (SBT) (Cartwright and Harary, 1956). SBT asserts that positive relations should exhibit transitivity (two positive paths should close with a positive edge) while negative relations follow anti-transitivity (two negative paths should close with a positive edge). Conversely, a positive two-path should not be closed by a negative relation, and a negative two-path should not be closed by another negative relation, as these patterns produce unbalanced triads.
@@
To explore these dynamics, our approach utilises the exponential random graph modelling (ERGM) framework (Lusher et al., 2013), which is particularly suited to modelling the complex interplay between positive and negative edge configurations. ERGMs allow us to incorporate terms that capture relational tendencies, such as the clustering of positive edges or the formation of unexpected negative edges between individuals with similar nodal covariate information. By accounting for these relational patterns, we can accurately assess the statistical significance of balanced configurations and their impact on the overall structure of the network (Fritz et al., 2025).
@@
In this paper, we examine undirected signed network panel data, which consists of cross-sectional snapshots of network relationships observed at regular intervals. In particular, we focus on the analysis of political relationships among U.S. Senators, measured across successive U.S. Congresses (Neal, 2014, 2020). Each snapshot captures the state of signed interactions within the Senate during each legislative year, offering insight into the evolving dynamics of political alliances and rivalries over time. We extend the temporal exponential random graph model (TERGM) framework (Robins and Pattison, 2001; Hanneke et al., 2010) to better accommodate the dynamics of signed interactions in longitudinal network data. Specifically, we incorporate a separable specification (Krivitsky and Handcock, 2014) that distinguishes between the formation and persistence of edges, enabling a more detailed understanding of temporal dependence. Building on the decomposition introduced by Lerner (2016), we further enhance the TERGM by modelling the signed network as the outcome of two interrelated processes: an interaction process and a conditional sign process. This extension allows for a more efficient and interpretable parameterisation that captures not only how relationships form and persist over time, but also how structural balance, through both balanced and imbalanced signed configurations, shapes the evolution of network interactions.
This perspective aligns with recent work showing that micro-level mechanisms can accumulate to produce pronounced patterns at the macro level (Duxbury, 2024). Even modest biases in the formation or sign of individual edges can have substantial effects on overall network structure, influencing cohesion, clustering, and balance (Amati et al., 2018). Incorporating endogenous micro-level effects into our separable framework ensures that both the local dynamics and their consequences for global network patterns are explicitly represented.
@@
The paper is structured as follows. In Sections 2 and 3, we provide an overview of ERGMs and temporal ERGMs respectively. Section 4 offers a comprehensive definition of signed networks and presents an overview of SBT. In Sections 5 and 6, we introduce a multi-layer interpretation of the signed process, distinguishing between the interaction process and the conditional sign process. We describe how these two components are incorporated into the ERGM generative framework to capture the dynamics of signed relationships. Section 7 combines temporal separability with layer separability to define 2-layer separable TERGMs. Here, we define a Markov process based on binary ERGM terms, which allows for a principled handling of Structural Balance Theory configurations. In Section 8, we outline a Bayesian inferential approach, which quantifies the uncertainty in the posterior model parameter estimates by using an adaptive approximate exchange algorithm (Haario et al., 2001; Caimo and Friel, 2011) to facilitate sampling from the doubly-intractable posterior distribution. In Section 9, we present a comparative simulation study to demonstrate parameter recoverability and the robustness of the multi-layer conditional approach, and to highlight the differences between this method, the TERGM approach of Fritz et al. (2025), and its separable temporal variant. Finally, Section 10 demonstrates the utility of our modelling approach by applying it to the analysis of political relationships among U.S. Senators. We assess model fit using posterior predictive checks to evaluate the model performance. Section 11 concludes with a discussion on potential extensions and future directions.
@@ 2 Exponential random graph models
where $\mathbf{s}(\mathbf{y})\in\mathbb{R}^{Q}_{>0}$ is a known vector of $Q$ sufficient statistics (Besag, 1974), $\boldsymbol{\varphi}\in\mathbb{R}^{Q}$ is the associated parameter vector, and $\kappa(\boldsymbol{\varphi})$ a normalising constant which is computationally intractable to evaluate for all but trivially small graphs. The dependence hypothesis at the basis of these models is that edges form small structures called configurations. There is a wide range of possible network configurations (Robins et al., 2007) which give flexibility to adapt to different contexts. A positive (or negative) value of the parameter $\varphi$ indicates a tendency for the configuration represented by $s(\mathbf{y})$ to appear more (or less) frequently in the data than would be expected under an Erdős–Rényi random graph model with an edge probability of 0.5.
@@ 3 Temporal exponential random graph models
Temporal exponential random graph models (TERGMs) (Robins and Pattison, 2001; Hanneke et al., 2010) describe the joint distribution of a dynamic network sequence by assuming a Markov process on the network from one time step to the next:
$p(\mathbf{y}_{1},\dots,\mathbf{y}_{T}\mid\mathbf{y}_{0},\boldsymbol{\varphi})=\prod_{t=1}^{T}p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\varphi}\right)=\prod_{t=1}^{T}\exp\{\boldsymbol{\varphi}^{\top}\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1}\right)-\kappa(\boldsymbol{\varphi};\mathbf{y}_{t-1})\},$
where $\boldsymbol{\varphi}$ parametrises the influence of sufficient network statistics $\mathbf{s}\left(\mathbf{y}_{t};\mathbf{y}_{t-1}\right)$ on the likelihood and $\kappa(\boldsymbol{\varphi},\mathbf{y}_{t-1})$ is a normalising constant. The process can be extended to incorporate higher-order Markov dependence by assuming that the network $\mathbf{y}_{t}$ depends on $K\in\{0,1,\cdots,T-1\}$ previously observed networks. This is achieved by including lagged networks in the network statistics $\mathbf{s}(\cdot)$ with $K$ selected appropriately (Leifeld et al., 2018).
@@
where $\boldsymbol{\varphi}^{\mathcal{F}}$ and $\boldsymbol{\varphi}^{\mathcal{P}}$ describe respectively the edge formation and persistence process from t1t-1 to $t.$ Then, given $\mathbf{y}^{\mathcal{F}},\mathbf{y}^{\mathcal{P}}$ and $\mathbf{y}_{t-1},$ we have $\mathbf{y}_{t}=\mathbf{y}^{\mathcal{P}}\cup\left(\mathbf{y}^{\mathcal{F}}\;\backslash\;\mathbf{y}_{t-1}\right).$
The STERGM framework addresses a key limitation of standard TERGMs by separating the modelling of edge formation and persistence. This separation allows for more interpretable modelling of dynamic networks, especially when we assume that edge formation and dissolution are governed by different mechanisms. For example, in a friendship network, a new friendship (formation) may be influenced by proximity while the maintenance of an existing friendship (persistence) may supported by the presence of mutual friends.
@@ 4 Signed network structures
A signed network consists of a set of $n$ nodes and signed edges between pairs of them. Its relational structure is described by a random adjacency matrix with elements of the set of edge variables $y_{ij}\in\{+1,-1,0\}$ respectively indicating a positive, negative or no relation between node $i$ and node $j.$ As for binary networks, signed networks may be directed or undirected depending on the nature of the relationships between the nodes. In this paper, we focus on signed undirected networks, although the methodologies introduced can be easily extended to the directed case.
Structural balance theory (SBT) examines how positive and negative relationships between actors evolve and ultimately stabilise within a network. It focuses on the dynamics and eventual equilibrium of these relationships in a signed network (Cartwright and Harary, 1956). SBT is commonly illustrated using the familiar concepts of friendship and enmity. These terms provide a clear and intuitive understanding of the fundamental principles of the theory. Balanced triadic configurations include: (1) a friend of a friend is a friend, and (2) an enemy of a friend is an enemy. In contrast, unbalanced configurations are: (3) a friend of an enemy is an enemy, and (4) an enemy of an enemy is a friend. Configuration (1) corresponds to positive triangles (Figure 1 (a)), while configuration (2) corresponds to triangles with two negative edges (Figure 1 (b)). These two patterns are part of the so-called strong balance structure (Harary, 1953). Davis (1967) later introduced the weak balance structure, under which configurations comprising triangles with all negative edges (Figure 1 (c)) are also regarded as balanced. Conversely, triadic configurations with exactly two positive edges and one negative edge (Figure 1 (d)) are considered unbalanced.
@@ 5 Multi-layer interpretation
Multi-layer ERGMs (Krivitsky et al., 2020) extend the traditional ERGM framework to model networks with multiple types of interactions between nodes, with each layer representing a distinct kind of relationship. Each layer is modelled using its own set of network statistics, capturing the structural tendencies specific to that type of interaction, while allowing for dependencies across layers.
For example, suppose $\mathbf{Y}_{1}$ represents a binary friendship network and $\mathbf{Y}_{2}$ represents a binary collaboration network among the same set of individuals. In this setting, any dyad can simultaneously exhibit both types of relations (e.g., two individuals may be friends, collaborators, or both). A multi-layer ERGM can include statistics such as transitivity or homophily within each layer (e.g., friends-of-friends in $\mathbf{Y}_{1}$, shared collaborators in $\mathbf{Y}_{2}$), while also incorporating cross-layer effects, such as whether an edge in $\mathbf{Y}_{1}$ increases the likelihood of an edge in $\mathbf{Y}_{2}$.
In the context of signed networks, where a pair of nodes cannot simultaneously share both a positive and a negative edge, the signed edge process can be decomposed as $\mathbf{Y}=\mathbf{X}\mathbf{Z},$ where $\mathbf{X},$ with dyadic elements $X_{ij}=1$ if an edge exists between nodes $i$ and $j$ and $X_{ij}=0$ otherwise, is a binary adjacency matrix encoding edge presence, and $\mathbf{Z},$ with dyadic elements $Z_{ij}=+1$ for positive edges and $Z_{ij}=-1$ for negative edges, encodes edge signs. This formulation ensures that each dyad has at most one edge with a single sign. Clearly, this can be viewed as a special case of a multi-layer network, where the sign information is meaningful only in the presence of an actual interaction. In other words, the sign layer is effectively constrained by the interaction layer. This dependency will be further discussed in the following section.
@@
where $\mathbf{s}(\mathbf{x},\mathbf{z})$ incorporates network statistics that model the two layers jointly (see Figure 2 (a)). The modelling approach for signed networks proposed by Fritz et al. (2025) is equivalent to a 2-layer ERGM which represents an ERGM with a uniform categorical reference measure where each dyad can take one of three possible states: positive, negative, or absent. The conditional probability that a dyad $(i,j)$ is positive, given the rest of the network $\mathbf{y}_{-(ij)}$ and the model parameter $\boldsymbol{\vartheta}$, is
$\Pr\left(Y_{ij}=+1\mid\mathbf{y}_{-(ij)},\boldsymbol{\vartheta}\right)=\frac{\exp\left\{\boldsymbol{\vartheta}^{\top}\mathbf{s}\left(\mathbf{y}_{ij}^{+1}\right)\right\}}{\sum_{y\in\{+1,-1,0\}}\exp\left\{\boldsymbol{\vartheta}^{\top}\mathbf{s}\left(\mathbf{y}_{ij}^{y}\right)\right\}},$
where $\mathbf{s}(\mathbf{y}_{ij}^{y})$ is the vector of network statistics evaluated when dyad $(i,j)$ is in state $y$.
@@ 6 Separable modelling
Lerner (2016) highlights the distinction between marginal and conditional tests of structural balance, showing that other network effects influencing the likelihood of interaction can mask the presence of balance. In such cases, structural balance may appear absent even when nodes actually prefer balanced triangles over unbalanced ones. To address this, a proper test of balance theory requires modelling the conditional probability of an edge being positive or negative, given that the edge exists. More recently, Lerner and Lomi (2020) have shown that, when conditioning on active users within a large relational event framework, balance theory provides a compelling explanation for the network structure of teams engaged in contentious tasks.
In line with this insight, our model explicitly separates the modelling of the interaction process ($\mathbf{X}$) from the sign process ($\mathbf{Z}$). Unlike the joint model described in the previous section, this separable approach allows the edge structure and edge signs to be modelled conditionally independently, while still accounting for their dependence through the conditional formulation. As shown by Lerner (2016) through analyses of multiple benchmark datasets commonly used in the literature, this separation offers both practical and interpretative advantages. It simplifies model estimation, and it reflects situations where the drivers of edge formation, such as affiliations, or spatial proximity, are distinct from those influencing edge polarity, such as trust, cooperation, or past conflict. For example, in a social network, whether an edge exists may be determined by organisational or logistical constraints, whereas the sign of that edge depends on the nature of the relationship. Formally, the marginal probability of the network $\mathbf{Y}$ can be decomposed as the product of the probability of interaction $\mathbf{X}$ and the conditional probability of the sign given interaction $\mathbf{Z}\mid\mathbf{X}$:
$\Pr(\mathbf{Y}=\mathbf{y})=\Pr(\mathbf{Z}=\mathbf{z}\mid\mathbf{X}=\mathbf{x})\;\Pr(\mathbf{X}=\mathbf{x}).$
The fact that $\mathbf{Z}$ is conditional on $\mathbf{X}$ does not imply a temporal ordering. The conditionality reflects the structural dependence of edge signs on the presence of edges and does not indicate that $\mathbf{Z}$ evolves after $\mathbf{X}$. This explicit separation clarifies how structural effects can be estimated for each process without conflating the mechanisms driving edge formation and edge sign. This leads to the following formulation of a separable ERGM:
@@
Figure 2: Graphical representation of (a) 2-layer ERGMs defined in Equation (3) and (b) separable 2-layer ERGMs defined in Equation (4). Squares represent observed variables (network layers); circles represent parameters and arrows represent directed dependencies. In panel (a), both $\mathbf{x}$ and $\mathbf{z}$ are influenced by a common parameter $\boldsymbol{\vartheta}.$ In panel (b), $\mathbf{x}$ and $\mathbf{z}$ each have their own parameters, denoted by $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$, respectively. These parameters are distinct, meaning the two processes are conditionally independent given $\mathbf{x}.$
7 Layered separable temporal models
The separable temporal assumption defined in Equation (2), when combined with the 2-layer separability assumption specified in Equation (4), results in a 2-layer STERGM, with the transition probability defined as:
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\right)=$ $\;p\left(\mathbf{y}^{\mathcal{F}}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{F}}\right)\times p\left(\mathbf{y}^{\mathcal{P}}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{P}},\boldsymbol{\zeta}^{\mathcal{P}}\right)$ (5)
$=$ $\;p\left(\mathbf{z}_{t}^{\mathcal{F}}\mid\mathbf{x}_{t}^{\mathcal{F}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{F}}\right)\times p\left(\mathbf{x}_{t}^{\mathcal{F}}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{F}}\right)\times$
$\;p\left(\mathbf{z}_{t}^{\mathcal{P}}\mid\mathbf{x}_{t}^{\mathcal{P}},\mathbf{y}_{t-1},\boldsymbol{\zeta}^{\mathcal{P}}\right)\times p\left(\mathbf{x}_{t}^{\mathcal{P}}\mid\mathbf{y}_{t-1},\boldsymbol{\xi}^{\mathcal{P}}\right)$
$=$ $\exp\left\{{\boldsymbol{\zeta}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{F}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{F}}};\mathbf{x}^{\mathcal{F}},\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\xi}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{F}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{F}}};\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\zeta}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{z}^{\mathcal{P}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\zeta}^{\mathcal{P}}};\mathbf{x}^{\mathcal{P}},\mathbf{y}_{t-1})\right\}\times$
$\exp\left\{{\boldsymbol{\xi}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{x}^{\mathcal{P}};\mathbf{y}_{t-1}\right)-\kappa({\boldsymbol{\xi}^{\mathcal{P}}};\mathbf{y}_{t-1})\right\},$